# A Neurosymbolic Framework for Interpretable Cognitive Attack Detection in Augmented Reality

### v1

@@ 4 Experiment
Table 1: Results of the state-of-the-art methods and CADAR on Extended AR-VIM Dataset
Method Attack Detection: Accuracy/F1-score (%) Localization Acc. of attack(%)
Text Visual Removal Addition Text+Visual Add.+Rem. Overall
CADAR (Ours) 80.2/ 74.9 82.3/ 76.1 83.0/ 86.3 82.1/ 83.6 67.0/ 69.7 86.1/ 75.2 80.1/77.6 68.0
GPT-o3 62.5/56.6 23.8/29.4 76.2/66.7 61.9/55.3 16.7/25.0 14.3/25.0 39.2/39.8 -
Gemini 2.5 66.7/71.8 66.7/77.8 81.0/69.4 90.5/88.4 14.3/18.1 71.4/71.4 69.4/71.1 -
3D ResNet 12.5/22.2 57.1/72.7 57.1/34.7 42.8/31.6 - - 42.4/40.3 -
ViViT 50.0/44.4 28.6/33.3 71.4/47.6 14.3/20.0 - - 41.1/36.3 -
In this section, we provide a detailed description of the CADAR implementation and the experimental settings used for our evaluation. We then present and analyze the numerical results of our experiments, demonstrating that CADAR significantly outperforms several baselines, achieving an overall accuracy of 80.1% compared to 69.4% for the next-best model. An ablation study is also conducted to analyze the contributions of our key components.
4.1 Implementation of CADAR
CADAR is implemented using a modular architecture. The perception graph is constructed and managed through several key components. The Node.py and Edge.py modules generate multimodal object representations by combining textual and visual features: GPT-4.1 is used for generating descriptive text, google/owlv2-base-patch16-ensemble (Minderer, Gritsenko, and Houlsby 2024) enables zero-shot object detection, and SAM2 (Ravi et al. 2024) provides object segmentation. Embeddings are extracted using all-MiniLM-L6-v2 for object names, all-mpnet-base-v2 for descriptions, and ResNet-34 for image features. The Graph.py module manages the perception graph structure, maintaining node–edge relationships and temporal information. Temporal reasoning is handled by NodeSet.py and EdgeSet.py, which update the graph across frames, detect potential attacks, and evolve the graph state. Attack detection is based on statistical anomaly detection using PCA to preserve 99% of variance, with thresholds computed from a 99% confidence level of the chi-squared distribution. Finally, Utils.py provides supporting tools for video processing and mathematical computations.
@@ 4.2 Experimental Settings
Extended AR-VIM Dataset.
We base our experiments on an extended version of the AR-VIM dataset, which is specifically designed to evaluate visual information manipulation in AR environments (Xiu and Gorlatova 2025). The Extended AR-VIM is the first publicly available AR cognitive attack dataset (details in Appendix). Each sample in the dataset consists of a short video collected by an AR device (Meta Quest 3 and Samsung Galaxy S24), approximately 5-10 seconds in length, where virtual content appears at a specific frame to trigger an attack. The dataset has 160 videos in different scenes, spanning all four cognitive-attack classes. The scene covers transportation (e.g., highway, train, aircraft), experimental settings (e.g., lab equipment), industrial environments (e.g., construction, agriculture), workspaces (e.g., warehouse, office), as well as medical, daily-life, and public-safety domains. These range from altering freeway exit-ramp signage to suppressing voltage warnings on live equipment. Together, these videos provide a diverse, realistic, and tightly controlled test bed for evaluating the perception-level of cognitive attacks in AR environments. The extended AR-VIM dataset consists of 41 removal attack samples, 35 addition attack samples, 40 text manipulation attack samples, 30 visual manipulation attack samples, and 14 mix-class attack (text + visual manipulation attacks and addition + removal attacks) samples. For baseline training models, it also contains 40 non-attacked videos.
Considering the scene availability, the dataset was collected using two complementary methods: (1) The monitor-based method places AR content on a monitor screen with background images and captures it via video recording, allowing rapid and controlled generation of attack instances; (2) The real-world-based method, in contrast, uses AR headsets to capture first-person views in physical environments, introducing natural visual noise and depth information. Employing both methods is critical to capturing the trade-off between efficient data generation and real-world fidelity.
Task and Metrics.
We evaluate our model on the task of attack classification in videos. CADAR supports frame-wise detection, variable-length multi-label outputs, and localization of multiple simultaneous attacks, which are capabilities that few existing models possess. As most baselines can only classify attacks at the video level and lack the capability for fine-grained, frame-level detection or localization, we compare CADAR against both advanced reasoning VLMs (GPT-o3 and Google Gemini 2.5 (Comanici et al. 2025)) and trained video classification models, including 3D-ResNet (Tran et al. 2018) and ViViT (Arnab et al. 2021). We use accuracy and F1-score to evaluate attack detection performance: frame-wise for CADAR, and video-level for the baselines. Additionally, we report attack localization accuracy, which measures how precisely the model locates the attacks in the frame. Results are summarized in Table 1. For mixed-attack instances, a detection is considered correct only if both attack types are correctly identified.
@@ 4.3 Numerical Results and discussion
Table 1 presents the evaluation results for attack classification on the extended AR-VIM dataset. CADAR outperforms both the supervised learning baselines and the pre-trained VLMs. Among the different attack types, removal and addition attacks are generally easier to detect, leading to higher accuracy compared to text and visual modifications. Specifically, CADAR achieves a detection accuracy of 83.0% for removal attacks and 82.1% for addition attacks, compared to 80.2% for text and 82.3% for visual attacks. Notably, CADAR demonstrates balanced and robust performance across all attack categories, achieving superior accuracy and F1 scores. For instance, its overall accuracy/F1 score is 80.1%/77.6%, significantly higher than the next best model, Gemini 2.5, which scores 69.4%/71.1%.
Supervised models such as 3D-ResNet and ViViT require large-scale training data and perform poorly in zero- or few-shot scenarios. For example, 3D-ResNet achieves an overall accuracy of 42.4%, and ViViT achieves 41.1%. In contrast, pre-trained VLMs leverage prior knowledge and exhibit strong reasoning abilities, but their outputs are often unstable and prone to hallucination. For example, GPT-o3 and Gemini 2.5 achieve overall accuracies of 39.2% and 69.4%, respectively.
CADAR offers a novel approach that addresses these limitations and achieves the best overall performance. However, it is not without shortcomings. The localization accuracy is constrained by the capabilities of the zero-shot object detection and segmentation models, which struggle in complex scenes or with unfamiliar objects. CADAR achieves a localization accuracy of 68.0%. Moreover, reduced performance on text attacks suggests that highly subtle edits, e.g., a single-character change that scarcely alters semantics or appearance, remain difficult for the detector.
@@ 4.4 Ablation Study
Figure 5: Ablation study on estimation module and reference frames.
Fig. 5(a) shows our ablation study examining the impact of the estimation module in particle filtering. When a polluted node that might have bad data due to VLM unstable output or due to an undetected attack, the estimation module’s weighting and resampling can effectively prevent it from corrupting the ground truth knowledge. In the ablation study, we choose 12 videos for all attack types. As shown in the left plot of the figure, for text and visual modification attacks, the detection accuracy drops significantly as the number of frames without the estimation module increases. Specifically, the accuracy for text attacks decreases from approximately $83\%$ to $69\%$ and from approximately $81\%$ to $55\%$ for visual attacks. This indicates that the estimation module is crucial for correctly identifying and mitigating errors in more subtle attack types. In contrast, for removal and addition attacks, the accuracy remains relatively stable, suggesting these attacks are more straightforward to detect and less dependent on the estimation module’s corrective capabilities.
@@
Fig. 5(b) shows the impact of the number of reference frames on detection accuracy. A larger number of reference frames, or more prior knowledge of the scene, will result in higher accuracy. For text and visual modification attacks, increasing the number of reference frames from 3 to 11 significantly improves performance. The accuracy for text attacks increases from around $55\%$ to $80\%$ and from around $50\%$ to $78\%$ for visual attacks. This demonstrates that a larger set of reference frames provides more prior knowledge for more accurate statistical reasoning, which is essential for detecting subtle text and visual changes. Similar to the estimation module study, the detection accuracy for removal and addition attacks remains largely unaffected by the number of reference frames, consistently staying above $80\%$. This reinforces the finding that these attacks are less sensitive to the amount of contextual information and prior knowledge.
@@ 5 Conclusion
We presented CADAR, a neuro-symbolic framework that couples a perception graph with particle-filter reasoning to yield interpretable attack detection in AR/MR. On the Extended AR-VIM dataset, CADAR performs strongly across all attack types, and ablations highlight the value of our estimation module and richer reference histories. Limitations stem from the underlying open-vocabulary detection/segmentation, and miss very fine-grained augmentations. Future work will pursue task-specific fine-tuning of perception models, and larger-scale real-world evaluations and user studies. Overall, CADAR highlights the value of neuro-symbolic design by combining the adaptability of foundation models with symbolic reasoning for a robust and interpretable AR safety mechanism.
Appendix A Extended AR-VIM Dataset Samples

### v3

@@ 4 Experiment
Table 1: Attack-wise Accuracy and F1-score (%) across scenes for frame-level and video-level baselines. Each attack type forms a block, with sub-rows indicating methods. Columns represent scene categories. CADAR outperforms VLM-based (GPT-5-mini, Gemini-2.5-flash) and supervised video models (3D-ResNet, ViViT) in 23 out of 25 attack–scene blocks.
Class-specific Accuracy / F1-score (%)
Attack Detection Method School&Office Traffic&Mobility Industrial&Infra Medical&Lab Public&Home Avg Acc/F1
[…]
GPT-5-mini 50.6 / 67.2 59.6 / 74.7 52.1 / 68.5 52.5 / 68.9 61.1 / 75.9 55.2 / 71.0
Gemini-2.5-flash 73.0 / 84.4 58.4 / 73.7 77.4 / 87.2 41.6 / 58.8 51.4 / 67.9 60.4 / 74.4
Video 3D ResNet 12.5 / 22.2 10.9 / 19.6 12.8 / 22.6 15.6 / 27.0 40.0 / 57.1 18.4 / 29.7
ViViT 26.3 / 41.7 70.6 / 82.8 25.0 / 42.1 33.3 / 50.0 43.5 / 60.6 39.7 / 55.4
[…]
GPT-5-mini 51.0 / 67.6 61.5 / 76.2 50.6 / 67.2 45.5 / 62.5 59.8 / 74.9 53.7 / 69.7
Gemini-2.5-flash 49.5 / 66.3 76.7 / 86.8 43.9 / 61.0 58.7 / 74.0 40.8 / 58.0 53.9 / 69.2
Video 3D ResNet 16.5 / 28.3 7.2 / 13.3 8.3 / 15.3 12.7 / 22.5 8.3 / 15.4 10.6 / 19.0
ViViT 22.2 / 36.4 44.4 / 61.5 20.0 / 33.3 28.6 / 44.4 25.0 / 40.0 28.0 / 43.1
[…]
GPT-5-mini 75.4 / 85.9 46.0 / 63.0 61.7 / 76.3 46.3 / 63.3 63.2 / 77.4 58.5 / 73.2
Gemini-2.5-flash 64.7 / 78.6 51.6 / 68.1 55.0 / 71.0 36.9 / 53.9 43.6 / 60.7 50.4 / 66.5
Video 3D ResNet 8.1 / 15.0 3.0 / 5.9 5.0 / 9.5 9.4 / 17.2 4.5 / 8.7 6.0 / 11.3
ViViT 50.0 / 66.7 54.5 / 70.6 46.2 / 63.2 71.4 / 83.3 41.7 / 42.3 52.8 / 65.2
[…]
GPT-5-mini 69.1 / 81.7 60.7 / 75.6 56.3 / 72.0 55.2 / 71.2 63.3 / 77.5 60.9 / 75.6
Gemini-2.5-flash 70.1 / 82.4 54.1 / 70.2 62.8 / 77.2 61.5 / 76.1 53.1 / 69.4 60.3 / 75.0
Video 3D ResNet 6.6 / 12.3 12.8 / 22.7 9.8 / 17.8 13.3 / 23.4 15.7 / 27.2 11.6 / 20.7
ViViT 30.0 / 46.2 45.5 / 62.5 20.0 / 33.3 25.0 / 40.0 55.6 / 71.4 35.2 / 50.7
[…]
GPT-5-mini 71.8 / 83.6 52.1 / 68.5 62.8 / 77.1 70.5 / 82.7 68.1 / 81.0 65.1 / 77.4
Gemini-2.5-flash 54.2 / 70.3 38.2 / 55.3 44.0 / 61.1 41.1 / 58.3 44.1 / 61.2 44.3 / 61.2
Video 3D ResNet 8.5 / 15.7 11.7 / 20.9 8.6 / 15.9 15.9 / 27.5 29.4 / 45.5 14.8 / 25.1
ViViT 40.0 / 57.1 15.4 / 26.7 7.7 / 14.3 40.0 / 57.1 30.8 / 47.1 26.8 / 40.5
Overall Mean (per method): CADAR: 74.6 / 84.2   GPT-5-mini: 58.7 / 73.4   Gemini-2.5-flash: 53.9 / 69.3   3D ResNet: 12.3 / 21.1   ViViT: 36.5 / 51.0
Table 2: Recall comparison for hybrid-attack scenarios(%), CADAR achieves the highest recall across both hybrid combinations.
Hybrid Attack Type Recall of Method (%)
GPT-5-mini Grok-4-fast CADAR Gemini-2.5-flash
Text+Visual 35.0 41.2 76.4 29.9
Obstruction+Injection 32.7 26.1 80.4 45.8
Table 3: Recall (%) of attack detection under different numbers of simultaneous attacks. CADAR maintains strong performance as the number of attacks increases, whereas baseline models degrade substantially.
Method Number of Attacks Avg
1 Attack 2 Attacks 3 Attacks 4 Attacks
GPT5-mini 66.4 40.2 31.5 29.6 41.9
Grok-4-fast 69.7 55.0 51.3 36.5 53.1
Gemini-2.5-flash 61.7 43.6 33.0 31.5 42.5
[…]
In this section, we provide a detailed description of the CADAR experimental settings used for performance evaluation. We then present and analyze the numerical results of our experiments, demonstrating that CADAR significantly outperforms several baselines. An ablation study is also conducted to analyze the contributions of our key components.
@@ 4.1 Experimental Settings
CADAR-50K Dataset.
The CADAR-50K dataset extends the AR-VIM dataset [11202433], which was designed to evaluate visual information manipulation in AR environments. CADAR-50K is the first publicly available dataset for AR cognitive-attack analysis, containing 375 videos across diverse scenes and more than 50K frames, covering all seven categories (Fig. 5 and Fig. 6). Each sample is an 8–12 second AR-captured video (recorded using a Meta Quest 3 or Samsung Galaxy S24), where virtual content is introduced at a specific frame to induce an attack. For the non-attack elements, we use irrelevant background objects, benign AR elements, and semantically benign prompts that can interfere with VLM outputs but do not alter the human-perceived task semantics.
Considering scene availability, the dataset was collected using two complementary methods: (1) a virtual-scene method, where fully synthetic environments are first generated and displayed on a monitor or VR headset, and AR content is then overlaid onto these virtual scenes before being captured through video recording. This enables rapid, controllable creation of diverse attack scenarios; and(2) a real-world method, where AR headsets record first-person views in physical environments, introducing natural noise, lighting variation, and depth cues.
Task, Baselines and Metrics.
Figure 5: CADAR-50K dataset overview.
We evaluate our model on the task of video-based cognitive attack classification. The CADAR dataset supports frame-wise detection, variable-length multi-label outputs, and localization of simultaneous attacks. We compare CADAR against two categories of baselines that similarly accept multiple reference frames as input and output attack classes: (1) multimodal VLMs, including state-of-the-art OpenAI GPT-5, Google Gemini-2.5 [comanici2025gemini] and xAI Grok-4, which perform fine-grained frame-level detection; and (2) supervised video models, the CNN-based 3D-ResNet [tran2018closer] and transformer-based ViViT [arnab2021vivit], which output a single prediction per video.
In our first experiment(Table 1), we use class-specific accuracy and F1-score to assess overall detection performance. The class-specific accuracy is defined as $\mathrm{TP}/(\mathrm{TP}+\mathrm{FN}+\mathrm{FP})$, which penalizes both missed detections and false alarms for that class while ignoring true negatives that dominate under severe class imbalance. For hybrid attack or multiple attack videos(Table 2 and 3), a prediction is considered correct only if all attack types are simultaneously identified. We also report per-class recall , $\mathrm{TP}/(\mathrm{TP}+\mathrm{FN})$, to measure correctness conditioned on the presence of the class.
@@ 4.2 Numerical Results and discussion
Cross-Scene Cognitive Attack Classification.
The results in Table 1 demonstrate that CADAR achieves the highest overall accuracy(74.6%) and F1-score(84.2%) and its average accuracy exceeding the next best baseline models (GPT-5-mini) by 16%. CADAR attains the highest class-specific accuracy and F1 in 23 out of 25 attack–scene blocks. The margin is especially large for Visual and Obstruction attacks, where CADAR delivers strong gains, underscoring its robustness to fine-grained semantic distortions. We also observe that VLM baselines exhibit substantial accuracy fluctuations across different scenes, indicating that their reasoning processes are unstable and prone to hallucination. Because they rely on jointly processing all textual and visual inputs, they become increasingly sensitive to misleading cues, malicious prompts, and complex visual environments. This vulnerability highlights the advantage of our neuro-symbolic approach, which offers more controlled and explainable reasoning under diverse AR scenarios.
On the other hand, traditional video-classification models (3D-ResNet and ViViT) perform poorly across all attack types, suggesting that the models are inadequate for capturing the fine-grained semantic inconsistencies central to cognitive attack detection. Their limited performance is further exacerbated by the lack of large-scale, AR-specific training data, making these models particularly weak in zero- or few-shot scenarios.
Overall, the results indicate that CADAR’s neurosymbolic design and temporal reasoning components provide strong advantages for detecting subtle, localized, and multi-instance AR cognitive attacks, achieving robust performance across diverse environments and attack conditions.
Figure 6: Visualization frame examples of CADAR outputs across five scene categories in CADAR-50K 375 videos. Rows show the Reference Scene, Detection Scene, and CADAR Output. Output attack types: I = Injection, V = Visual, T = Text, O = Obstruction, N = Non-attack. Some examples contain hybrid or multiple attacks.
Hybrid and Multiple Attacks
Table 2 and 3 evaluate model robustness in hybrid and multi-attack settings. In hybrid-attack scenarios, CADAR attains the highest recall for both Text+Visual(76.4%) and Obstruction+Injection(80.4%) cases and surpasses all VLM baselines by a substantial margin. This demonstrates that the proposed neurosymbolic graph reasoning framework can effectively detect the attacks separately. In multiple attack scenarios, the performance of VLM baselines declines sharply as attack complexity grows, whereas CADAR retains strong recall, from 81.0% for a single attack to 67.2% for four simultaneous attacks, showing the smallest performance drop among all compared methods. These results highlight CADAR’s resilience in multi-attack scenarios and its clear advantage in detecting objects in parallel.
Effect of Scene Complexity.
The recall trends in Fig. 7(a) reveal a clear divergence between VLM-based baselines and CADAR as scene complexity increases. For GPT-5-mini and Gemini-2.5-flash, recall drops sharply when the number of key objects and frames increases, suggesting that conventional VLMs struggle with reasoning over long visual and textual context and are easily affected by irrelevant or distracting content. In contrast, CADAR performance remains stable as reference frames increase, since it extracts complex scene information into a structural spatial-temporal perception graph. CADAR improves its decision accuracy by accumulating prior evidence through additional reference frames, and evaluates objects and attacks separately rather than relying on a single holistic reasoning step. The CADAR design results in substantially smaller performance degradation compared with standard VLMs.
@@ 4.3 Ablation Study
Figure 7: Recall under increasing reference frames and object quantities (a), and ablation studies on the estimation module (b) and the number of reference frames (c).
Fig. 7(b) shows our ablation study examining the impact of the estimation module in particle filtering. When a polluted node that might have bad data due to VLM unstable output or due to an undetected attack, the estimation module’s weighting and resampling can effectively prevent it from corrupting the ground truth knowledge. We randomly sample 6 videos per base attack type (24 in total) from the test split for the ablation study. As shown in the left plot of the figure, for text and visual modification attacks, the recall drops significantly as the number of frames without the estimation module increases. Specifically, the recall for text attacks decreases from approximately $81\%$ to $63\%$ and from approximately $79\%$ to $55\%$ for visual attacks. This indicates that the estimation module is crucial for correctly identifying and mitigating errors in more subtle attack types. In contrast, for obstruction and injection attacks, the recall remains relatively stable, suggesting these attacks are more straightforward to detect and less dependent on the estimation module’s corrective capabilities.
@@
Fig. 7(c) shows the impact of the number of reference frames on recall. A larger number of reference frames, or more prior knowledge of the scene, will result in higher recall. For text and visual modification attacks, increasing the number of reference frames from 3 to 11 significantly improves performance. The recall for text attacks increases from around $52\%$ to $83\%$ and from around $48\%$ to $79\%$ for visual attacks. This demonstrates that a larger set of reference frames provides more prior knowledge for more accurate statistical reasoning, which is essential for detecting subtle text and visual changes. Similar to the estimation module study, the recall for obstruction and injection attacks remains largely unaffected by the number of reference frames, consistently staying around $80\%$. This reinforces the finding that these attacks are less sensitive to the amount of contextual information and prior knowledge.
@@ 5 Conclusion
We presented CADAR, a neuro-symbolic framework that couples a perception graph with particle-filter reasoning to yield interpretable attack detection in AR/MR. On the CADAR-50K dataset, CADAR performs strongly across all attack types, and ablations highlight the value of our estimation module and richer reference histories. Overall, CADAR highlights the value of neuro-symbolic design by combining the adaptability of foundation models with symbolic reasoning for a robust and interpretable AR safety mechanism.
Appendix
Appendix A CADAR Implementation Details
This section provides the full implementation details of CADAR, covering perception-graph construction, multimodal embedding generation, temporal reasoning, and statistical attack detection. The system follows a modular architecture implemented across the Node.py, Edge.py, Graph.py, NodeSet.py, EdgeSet.py, and Utils.py modules in our codebase.
A.1 Perception Graph Construction
Object Detection and Segmentation.
For each frame, CADAR extracts object proposals using two vision modules:
  • Zero-shot object detection:
    google/owlv2-base-patch16-ensemble [minderer2024scalingopenvocabularyobjectdetection], which produces category-agnostic bounding boxes and confidence scores.
  • Segmentation masks: SAM2 [ravi2024sam2segmentimages], used to obtain object-level silhouettes for robust visual embedding extraction.
Textual Attribute Generation.
To obtain semantic descriptions associated with the detected region, we use a lightweight VLM (GPT-4.1) generates short, structured textual descriptions. These descriptions include attributes such as object type, color, role, contextual relevance, and scene relationships. The attributes also include numerical importance, reasonability and probability.
Multimodal Embedding Extraction.
Each detected object is represented by three embedding modalities, each extracted using a specialized encoder:
  • Name embedding: paraphrase-MiniLM-L6-v2 (384-d).
  • Description embedding: all-mpnet-base-v2 (768-d).
  • Visual embedding: a ResNet-34 feature encoder applied to SAM2-extracted object crops (512-d).
These multimodal perceptual vectors are then projected using PCA for statistical reasoning (described below).
Node and Edge Formation.
Implementation Details.
The software architecture is structured around two primary classes, Node.py and Edge.py, which manage the lifecycle and serialization of the perception graph components. The Node class orchestrates the bottom-up extraction pipeline, initializing the vision models to generate object masks and invoking the VLM to populate attribute fields (e.g., reasonability, importance) while computing and storing dense multimodal embeddings. Complementing this, the Edge class executes the relational analysis by ingesting the instantiated node list and source image; it prompts the LLM to infer pairwise dependencies (such as spatial or functional links) and encodes these descriptions into vector format, thereby establishing the graph’s connectivity before serializing the complete state to compressed binary files.
A.2 Graph Backbone and Temporal Tracking
Graph Management.
The Graph.py module serves as the fundamental container for encapsulating the state of the perception system at discrete timestamps. It aggregates detected entities and their pairwise relationships into a coherent network structure, storing lists of Node and Edge instances alongside temporal metadata and references to the source imagery. Beyond acting as a static data structure, this module facilitates the serialization of the complete graph state into compressed binary formats to ensure efficient persistence and provides visualization utilities that render the probabilistic attributes and topological connectivity of the scene using network layout algorithms.
NodeSet and EdgeSet Evolution.
To transform isolated frame-level detections into continuous historical tracks, the system relies on the NodeSet.py and EdgeSet.py modules. The NodeSet class maintains the identity of objects over time by associating new detections with existing tracks via multi-modal embedding similarity (visual, semantic, and nominal) rather than simple spatial overlap. It implements a particle filter-based mechanism to estimate stable node attributes from noisy observations and employs statistical methods—such as PCA and Mahalanobis distance—to detect anomalies or adversarial attacks. Complementing this, the EdgeSet tracks the evolution of semantic and spatial relationships between these persistent entities, preserving the history of interactions to support high-level reasoning as the scene evolves.
A.3 Statistical Anomaly Detection
Dimensionality Reduction and Feature Whitening.
To ensure robust distance metrics in high-dimensional spaces, we apply Principal Component Analysis (PCA) to the embedding history of each track. For a set of historical embedding vectors $\mathbf{X}\in\mathbb{R}^{N\times D}$, we compute a projection that preserves a specific variance ratio. Specifically, for node textual and visual embeddings, we retain 99% of the variance ($v=0.99$), while for edge description embeddings, we retain 90% ($v=0.90$). For visual embeddings, we additionally apply whitening to normalize the principal components to unit variance. This projection results in a dynamic latent dimension $k$ (typically $16\leq k\leq 32$), which defines the degrees of freedom for subsequent statistical tests.
Regularized Mahalanobis Distance.
We model the distribution of ”clean” reference frames for any given object or relationship as a multivariate Gaussian. To assess whether a new observation $x_{t}$ belongs to this distribution, we calculate the squared Mahalanobis distance. To prevent numerical instability when the covariance matrix is singular (common in short history windows), we introduce a regularization term $\epsilon\mathbf{I}$:
$d^{2}(x_{t})=(x_{t}-\mu)^{\top}(\Sigma+\epsilon\mathbf{I})^{-1}(x_{t}-\mu)$ (22)
where $\mu$ is the mean of the historical embeddings, $\Sigma$ is the covariance matrix, and $\epsilon=10^{-6}$ is a smoothing factor.
Chi-Square Hypothesis Testing.
Anomalies are identified by comparing the squared distance against a critical value derived from the Chi-square distribution with $k$ degrees of freedom. An observation is flagged as an attack (e.g., semantic modification or visual spoofing) if:
$d^{2}(x_{t})>\chi^{2}_{1-\alpha,\,k}$ (23)
where $\alpha$ represents the significance level. We employ adaptive sensitivity thresholds: $\alpha=0.10$ for textual deviations, $\alpha=0.05$ for visual deviations, and $\alpha=0.01$ for edge relationship deviations. This statistical framework allows CADAR to detect subtle adversarial manipulations that deviate from the established semantic baseline.
A.4 Particle-Filter-Based Temporal Reasoning
To mitigate the stochasticity of VLM outputs and detection noise, we implement a particle filter that maintains a belief state over node attributes. Unlike standard tracking which estimates position, our filter estimates semantic consistency.
Weight Update.
For a node $n$ with history $H$, we define the weight $w_{i}$ of the $i$-th particle (attribute hypothesis) based on its similarity to the incoming observation $z_{t}$. For scalar attributes (probability, importance, reasonability), weights follow a Gaussian kernel:
$w_{i}\propto\exp\left(-\frac{(v_{i}-z_{t})^{2}}{2\sigma^{2}}\right)$ (24)
For high-dimensional embedding attributes, weights are derived from cosine similarity, scaled by a confidence factor $\lambda$:
$w_{i}\propto\exp\left(\lambda\cdot\frac{\mathbf{e}_{i}\cdot\mathbf{z}_{t}}{\|\mathbf{e}_{i}\|\|\mathbf{z}_{t}\|}\right)$ (25)
where $\lambda=1$ is the scaling factor used to tune the confidence in the top matches.
Resampling and Estimation.
We perform importance resampling to filter outlier descriptions generated by the VLM. The final estimated state $\hat{x}_{t}$ is computed as the weighted centroid of the resampled population, effectively smoothing the semantic trajectory of the object and rejecting transient classification errors.
Table 4: Hyper-parameters for Statistical Detection and Tracking.
Parameter Value Description
Matching Threshold 0.2 Similarity floor for node association
Min. History Window 5 frames Minimum samples required for PCA/Mahalanobis
PCA Variance (Nodes) 0.99 Variance retained for text/image embeddings
PCA Variance (Edges) 0.90 Variance retained for edge descriptions
Regularization ($\epsilon$) $10^{-6}$ Covariance matrix stabilization term
$\alpha$ (Text) 0.10 Significance level for node text anomalies
$\alpha$ (Image) 0.05 Significance level for visual appearance anomalies
$\alpha$ (Edge) 0.01 Significance level for relationship anomalies
Particle Scale ($\lambda$) 1.0 Exponential scaling for embedding similarity weights
A.5 Supporting Utilities
Utility and Pre-processing.
The utils.py module serves as a comprehensive toolkit for low-level visual processing and mathematical operations underpinning the higher-level graph logic.
A.6 Summary
Overall, CADAR integrates zero-shot perception, multimodal symbolic representation, PCA-based statistical modeling, and particle-filter temporal reasoning into a unified framework for robust AR cognitive attack detection. The modular architecture allows flexible substitution of detectors, encoders, and reasoning mechanisms, while ensuring reproducibility across all experimental settings.
Appendix B CADAR-50K Dataset Samples